OpenStudy (vgonzalez):

The function f(t) = 4 cos(pi over 3t) + 15 represents the tide in Bright Sea. It has a maximum of 19 feet when time (t) is 0 and a minimum of 11 feet. The sea repeats this cycle every 6 hours. After five hours, how high is the tide? 13.5 feet 16 feet 18.5 feet 17 feet

OpenStudy (vgonzalez):

@elementwielder

OpenStudy (campbell_st):

well substitute t = 5 and see what happens

OpenStudy (vgonzalez):

how would you do that plug in the numbers into formula?

OpenStudy (anonymous):

sorry idk

OpenStudy (vgonzalez):

okay so would i did cos(pie/3 x 5) and got 0.999

OpenStudy (campbell_st):

well its \[f(5) = 4 \times \cos(\frac{\pi}{3} \times 5) + 15\] is your calculator set for radians or degrees...?

OpenStudy (vgonzalez):

i think radians. not sure. I'm trying to solve the pie/3 x 5) first and got a different answer.. pie/243.

OpenStudy (campbell_st):

well try (5 times pi)/3

OpenStudy (vgonzalez):

5.23?

OpenStudy (campbell_st):

no you're not trying to solve it... you just need to evaluate it... when t = 5

OpenStudy (campbell_st):

ok... so work in degrees..... pi/3 = 60 so use \[f(5) = 4\times \cos(60 \times 5) + 15\]

OpenStudy (vgonzalez):

okay so far i did the 60 x 5 = 300 Now i just need to find cosine next correct?

OpenStudy (campbell_st):

yep... thats right

OpenStudy (campbell_st):

makes sense to me... and if you had the calculator in radians, you would have the same answer